A 200 kg load is hung on a wire having a length of 4 m, cross-sectional area 1.2 x10-4 m2 , and Young‘s modulus 8 x 1010 N/m2 . What is its increase in length?


1
Expert's answer
2021-11-15T13:04:58-0500

"m=200 \; kg \\
L=4 \; m \\
Y = 8 \times 10^{10} \; N/m^2 \\
A = 1.2 \times 10^{-4} \; m^2 \\
Y = \frac{f/A}{ΔL/L} \\
ΔL = \frac{200 \times 9.8 \times 4}{1.2 \times 10^{-4} \times 8 \times 10^{10}} \\
ΔL = 0.816 \times 10^{-3} \; m"